1(a) a vessel containing an unknown volume of argon gas in cm3 at temperature 300f and pressure 101kpa was inverted and placed in cold ethanol. As the gas contracted, the pressure was quadrupled while the temperature reduced by one-third. What is the final volume of gas in dm3?
(B.) Calculate the root mean square of bromine molecule at temperature 23 degree Celsius and at pressure 1.00 atm. What is the root mean square of the same gas at temperature 23 degree Celsius and pressure 1.5 atm
1
Expert's answer
2019-05-02T07:30:55-0400

(a) PV = nRT

T1 = 300 0F = (5/9 x (300 - 32)) + 273 = 422 K;

T2 = 300/3 = 100 0F = 311 K

P1 = 101 kPa = 101000 Pa

P2 = 101 kPa x 4 = 404 kPa = 404000 Pa

P1V1 / T1 = P2V2 / T2

101000 x V1 / 422 = 404000 x V2 / 311

V2 = 0.001 x V1

 

(b) Urms = (3p/ρ​)1/2

1.​ ρ = (p x M) / (R x T);

M(Br2) = 2 x 80 = 160 g/mol;

R = 0.082 (atm x L)/(mol x K);

T = 23 + 273 = 296 K

ρ = (1 x 160) / (0.082 x 296) = 6.59 g/L;

Urms = (3 x 1 / 6.59)1/2​ = 0.4554 m/s

2. ρ = (1.5 x 160) / (0.082 x 296) = 9.89 g/L;

Urms = (3 x 1,5 / 9.89)1/2​ = 0.4555 m/s.